Determine the inverse z-transform of X(Z)=e^z, ROC:all z except |z|=infinity


1
Expert's answer
2022-02-07T20:46:02-0500

X(Z)=e^z=1+Z/1+ Z2/2+Z3/6..........................

Inverse z-transform

X(n)=(.................1/6+1/2+1+1)